OpenStudy (anonymous):

whats the volume of the solid obtained when rotationg the region under teh graph of the function f(x)= 2/ (x+1) on the interval [0,1].. NEED HELP

OpenStudy (anonymous):

go to the earlier one creeg

OpenStudy (anonymous):

ok i'll repeat it here u there ?

OpenStudy (anonymous):

volume = INT pi* 4/(x+1)^2 dx btw 0 and 1 = pi( 4 * (x+1)^-1 / -1) btw 0 and 1 = pi (4*(2)^-1 /-1 - (4*1^-1 / -1)) = pi(-2+4) =2pi

OpenStudy (anonymous):

ok ? thats a bit easier computation compared to the last one.

OpenStudy (anonymous):

the previous one - i messed up the square of x^(13/12) = x^(13/6) - not what i got it. but the method was right.

OpenStudy (anonymous):

just in case u r confused - at first i thought this post was the same as ur first one,